The wing of an airplane has an average cross-sectional area of 13 m2 and experiences a lift force of 78,000 N. What is the average difference in the air pressure between the top and bottom of the wing?
1
Expert's answer
2019-04-17T12:12:01-0400

The lift force

"F=\Delta P\times A"

So, the average difference in the air pressure between the top and bottom of the wing

"\Delta P=\frac{F}{A}=\frac{78\:000\:\rm{N}}{13\:\rm{m^2}}=6000\:\rm{Pa}=6\:\rm{kPa}"